13

Suppose there are three buyers of candy in a market: Tex, Dex, and Rex. The market demand and the individual demands of Tex, Dex, and Rex are shown in the table below. a. Fill in the table (gray-shaded cells) for the missing values. Individual Quantities Demanded Price per Candy Total Quantity Demanded Tex Dex Rex S8 13 20 27 4 15 19

b. Which buyer demands the least at a price of $5? Tex The most at a price of $7? (Click to select) c. Which buyers quantity demanded increases the most when the price is lowered from $7 to $6? (Click to select) d. Which direction would the market demand curve shift if Tex withdrew from the market? (Click to select) What if Dex doubled his purchases at each possible price? Click to select) e. Suppose that at a price of $6, the total quantity demanded increases from 20 to 30 . Is this a “change in the quantity demanded or a change in demand?
Answer #1

Answer:

Question A)

Since Total Quantity demanded = Tex Quantity+Dex Quantity+Rex Quantity

Price Tex Dex Rex Total Quantity demanded
8 1 2 3 6
7 3 3 7 13
6 5 4 11 20
5 7 5 15 27
4 9 6 19 34

Question B)

Dex will demand the least at $5.

Rex will demand the Most at $7.

Question C)

If price changes form$7 to $6 then

Increase in demand for Tex=5-3=2

Increase in demand for Dex=4-3=1

Increase in demand for Rex=11-7=4

from above calculation it is clear that Increase in demand for Rex is the most.

Question D)

If Tex withdraw from the market then new quantity demanded is

Price Total Quantity demanded New Total Quantity demanded
8 6 5
7 13 10
6 20 15
5 27 20
4 34 25

From above table it is clear that for a particular price there is decrease in demand. So demand curve will shift to left.

Price $6-__ D1 15 20 Quantity

if Dex doubled his purchase

Price Total Quantity demanded New Total Quantity demanded
8 6 8
7 13 16
6 20 24
5 27 32
4 34 40

From above table it is clear that for a particular price there is increase in demand. So demand curve will shift to right.

Price $6--_ D1 20 Quantity

Question E)

since there is increase in demand for a particular price so there is change in demand. So there is shift in demand curve towards right.